OpenStudy (anonymous):

Two 6-sided dice are rolled. What is the probability the sum of the two numbers on the dice will be 9? Answer 8/36 1/9 9/10 9

OpenStudy (anonymous):

Number of outcomes= 6^2 = 36 Combinations that equal 9: 3+6, 4+5, 5+4, 6+3 =4 outcomes 4/36 = 1/9

OpenStudy (anonymous):

There are four ways to roll nine (3,6), (4,5), (5,4), (6,3) out of the thirty six possible ways the dice might land. That gives a probability of about one of nine.

OpenStudy (anonymous):

I clearly need to work on my typing speed...LOL

OpenStudy (anonymous):

LOl